# Deep Probabilistic Direction Prediction in 3D with Applications to Directional Dark Matter Detectors

### v1

@@ 1 Introduction
In order to clarify our contribution, we summarize related work in three related areas. In Section 2.3, we discuss the use of directional statistics for learning deep embedding spaces. In Section 2.1, we review the literature on deep models for predicting direction using non-probabilistic methods or methods that do not use 3D geometry. In Section 2.2, we review the literature for the related problem of probabilistically predicting 3D orientations.
@@
Through experiments we show that our probabilistic direction predictor significantly outperforms traditional algorithms when applied to a simulated electron recoil data set. Leveraging our framework’s uncertainty predictions, performance can be enhanced by discarding high-uncertainty samples. By discarding only the top 1% of events with highest directional uncertainty, our model achieves the best expected directional performance (as defined in A). This development is far-reaching as it has the potential to enhance the sensitivity of all directional recoil detection experiments. Exciting possibilities include aiding in the discovery of dark matter (Billard et al., 2014; O’Hare, 2021) or resolving the solar abundance problem (Villante and Serenelli, 2019; O’Hare et al., 2022). The framework can also be applied to general problems where direction needs to be predicted, we demonstrate this by applying it to the task of detecting the directions 3D arrows in B.
Our contributions include: (1) The first deep learning framework that predicts 3D direction probabilistically, (2) a derivation of the negative log likelihood loss function and approximations that stabilize training, (3) two experiments that empirically demonstrate learning and calibration, something which is lacking in related work, and (4) a deep model for predicting the direction of electron recoils from 3D point cloud data using sparse 3D convolutional networks.
@@ 2.3 Deep embedding models
Hyperspheres have long been popular manifolds for embedding data. Embedding data on the curved hypersphere provides a natural way to balance the competing attraction of similar data and repellence of dissimilar data. For example,  Elad et al. (2005) use multidimensional scaling to embed data on $\mathbb{S}^{2}$. For the same reason, many deep learning models constrain latent space representations to a hypersphere, for example in variational autoencoders Davidson et al. (2018) or in contrastive learning He et al. (2020); Chen et al. (2020). Some of this work makes use of the vMF distribution to quantify the variance in heteroscedastic predictions on the hypersphere latent space Davidson et al. (2018). However, these embedding methods are unsupervised in that the embeddings on the hypersphere are learned to optimize some objective. This work concerns the supervised learning scenario in which fixed targets live on the sphere.
@@ 3.2 Practical considerations
Some considerations are required to make the likelihood function and gradients of the probabilistic model tractable. First, the $f(\kappa_{i})\equiv\ln{\frac{\kappa_{i}}{4\pi\sinh{(\kappa_{i})}}}$ term on the right-hand side of Equation 4 is numerically unstable. When $\kappa_{i}=0$, the vMF distribution in Equation 3 is uniform over the sphere, but during learning $\kappa_{i}$ generally increases as the distribution becomes more localized (uncertainty decreases). If $\kappa_{i}\geq 87$, the denominator of $f(\kappa_{i})$ evaluates to zero (assuming a 32-bit floating point data type) resulting in an infinite loss. To avoid this, we use the limit
$\lim_{\kappa_{i}\to\infty}f(\kappa_{i})=\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}.$
In fact, $f(\kappa_{i})\approx\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}$ is a well-known approximation even for small values of $\kappa_{i}$ Dhillon and Sra (2003). For example, if $\kappa=9$, the values of $\ln(\frac{\kappa_{i}}{4\pi\sinh{(\kappa_{i})}})$ and $\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}$ are indistinguishable when using the 32-bit floating point data type. Therefore, we replace the loss definition from Equation 4 with
$\textrm{NLL}=\frac{-1}{N}\sum_{i=1}^{N}\mathbbm{1}_{\kappa_{i}<9}\left(f(\kappa_{i})\right)+\mathbbm{1}_{\kappa_{\geq}9}\left(\ln\left(\frac{\kappa_{i}}{2\pi}\right)-\kappa_{i}\right)+\kappa_{i}(\mathbf{y}_{i}\cdot\hat{\mathbf{y}}_{i}).$ (5)
This function can be implemented in PyTorch (Paszke et al., 2019) using the torch.where function, but unfortunately at the time of this writing, PyTorch produces a nan gradient if either of the terms evaluates to infinity (Paszke et al., 2019). Thus we solve this problem by removing the term $f(\kappa_{i})$ from Equation 5 completely, approximating it with its 15th-order Taylor Series about $\kappa_{i}=0$
$T_{15}(\kappa_{i})\triangleq\sum^{15}_{0}\frac{f^{(n)}(0)}{n!}\kappa_{i}^{n}=\ln(4\pi)+\frac{\kappa^{2}}{6}-\frac{\kappa^{4}}{180}+\frac{\kappa^{6}}{2835}-\frac{\kappa^{8}}{37800}+\dots.$
@@
The condition $\kappa_{i}<2.65$ is chosen so that Equations 4 and 6 agree within a margin of $0.14\%$.
@@ 4 Application to Electron Recoils
For a demonstration of our method on a simple toy problem see B. Here, we apply it to directional dark matter detectors. The most mature directional recoil detection technology is the gas time projection chamber (TPC). Gas TPCs are capable of determining 3D direction of recoiling atomic nuclei or electrons by reconstructing their ionization tracks. The detectors produces a 3D voxel grid, where each voxel provides a measurement of ionization at that position, as illustrated in Figure 1. We focus on determining the initial direction of the electron recoil signal because it represents the most complex scenario. There is also abundant literature on a non-machine learning algorithm for determining the direction of electrons in gas detectors (Weisskopf et al., 2016; Marco et al., 2022) which serves as a well-studied non-machine learning baseline in our experiments.
@@ 4.1 Simulation, Preprocessing, and Data splits
The final dataset contains 2,766,798 simulations. A random data split is used to create a training set (80%) and a validation set (20%). To create the test data sets we simulate another $2\times 10^{4}$ electron recoils at $40$ and 505050\,50keV. We follow the same steps outlined above except for a slight modification. The applied Gaussian smearing is no longer drawn from a uniform distribution. Instead, we probe two specific cases: a high diffusion case where $443\,{\upmu}{\rm m}$ of smearing is applied and a low diffusion case where $200\,{\upmu}{\rm m}$ is applied. The two energies and two diffusion cases give us four test data sets.
@@ 4.2 Architecture
The output of the 3D convolution layers is flattened and then passed to a sequence of densely-connected layers that produce the parameters of the vMF distribution (Figure 4b). One arm ends in three linear neurons with L2-normalization, interpreted as the predicted direction on $\mathbb{S}^{2}$ ($\hat{\mathbf{y}}_{i}$), while a second arm ends with a single output neuron with a Softplus activation function, interpreted as the predicted uncertainty ($\kappa_{i}$). To evaluate the advantage of the probabilistic approach, we also train a non-probabilistic version in which the uncertainty arm is removed (Figure 4a). We refer to the probabilistic, heteroscedastic convolutional network as HCN and this non-probabilistic, “regular” convolutional network as RCN. Further details of the dense portions of our models are in Figure 4.
Figure 4: The dense portion of our models. The feature extraction described in Figure 3 is followed by either the left or the right neural network to make the RCN or HCN model, respectively. The illustrations do not display the actual number of neurons per layer. The number of neurons and the activation functions are specified in the legend. A learnable bias is added to every layer.
@@ 4.3 Training
The RCN and HCN models are trained by minimising the loss functions outlined in Equations 1 and 6, respectively, using the Adam optimizer (Kingma and Ba, 2014). The learning rate is set to $\textrm{lr}=0.0001$, the betas values are $(0.94,0.999)$, and $\textrm{eps}=10^{-7}$, while all other parameters are kept at the PyTorch default value. The deep learning models are trained with mini-batches of size $N=256$. Hyperparameters were optimized by exploring values in the range $\textrm{eps}\in[10^{-8},10^{-6}]$, $\textrm{lr}\in[10^{-6},10^{-4}]$, and $N\in[64,256]$.
Both the RCN and HCN models are trained on the training set while their loss on the validation set is used for early stopping and hyperparameter selection. If the validation loss has not decreased in the last five epochs, training is stopped. The weights corresponding to the lowest validation loss are saved and used for the final model.
@@ 4.4 Performance
We evaluate the trained HCN model on the independent test sets discussed in Section 4.1 by comparing it to a commonly used non-machine learning algorithm and to our estimation of the best expected directional performance. The non-machine learning algorithm (Non-ML) is adapted from Marco et al. (2022). We generalize the original algorithm to 3D and explore different variations where we tune its parameters on the test sets and give the algorithm the correct head-tail. The best expected directional performance (Best-Expected) is estimated by utilizing information that is only available in simulation. The purpose of Best-Expected is not to be a viable option for analysis but to provide a benchmark of exceptional performance. Non-ML and Best-Expected are detailed in A. Furthermore, the RCN model provides a deep learning baseline to compare HCN to.
@@
To compare all of our models, we use the cosine distance loss in Equation 1. In A we explain that the Non-ML algorithm has a track efficiency based on the parameters used. Hence, when comparing the models we plot the cosine distance loss versus the percentage of omitted tracks, referred to as the efficiency cut. In Figure 5, we display the performance of all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing (left) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing (right). In both cases, we find that the deep learning models significantly outperform Non-ML, even under the most optimistic assumptions. On the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing and at a $0\%$ efficiency cut, the RCN, HCN, and Best-Expected models achieve a cosine distance loss of $0.104$, $0.104$, and $0.098$, respectively. On the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing and at a $0\%$ efficiency cut, the models achieve a cosine distance loss of $0.0624$, $0.0632$, and $0.0569$, in the same order. The HCN and RCN models have similar performance. However, the accuracy of the HCN model can be improved by discarding events with low $\kappa_{i}$ (high predicted uncertainty). The HCN model outperforms Best-Expected with just a $1\%$ efficiency cut, in both cases.
@@
Figure 5: Cosine distance loss versus efficiency cuts for all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing (left) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing (right). The RCN (red diamond) and Best-Expected (horizontal line) models have no efficiency cuts. For the HCN model, we improve accuracy by excluding examples with higher predicted uncertainty, results are presented with a blue curve. The dashed vertical line indicates when the HCN model beats Best-Expected. The orange point and black star indicate the performance of the Non-ML method with standard parameters and with parameters tuned on the test set, respectively. The downward arrows indicate how much improvement is possible if we help the Non-ML method by providing it with the true head-tail.
@@
To assess the calibration of the HCN model we plot a 2D histogram of the predicted uncertainty ($\kappa_{i}$) and the angle from the true direction to the predicted direction ($\theta_{i}$) for all test data, displayed in Figure 5(a). The spread in $\theta_{i}$ gets smaller for higher values of $\kappa$, indicating that the HCN model is predicting $\kappa_{i}$ appropriately. The solid red curve in Figure 5(a) shows the mean angle, calculated for each bin in $\kappa$. Using Equation 3, we can also calculate the expected mean angle as a function of $\kappa$
@@
where $I_{o}$ is a modified Bessel function of the first kind. The expected mean angle as a function of $\kappa$ is plotted as the dashed black line in Figure 5(a). The agreement of the expected mean angle with the mean angle for each bin in kappa suggests that the HCN model is well-calibrated.
@@
Building onto the 2D histogram in Figure 5(a), we check whether the predicted directions in a $\kappa$ bin behave as if they are drawn from a von Mises-Fisher distribution with that value of $\kappa$. For each bin in $\kappa$, we compute a distribution of $\mathbf{y}_{i}\cdot\hat{\mathbf{y}}_{i}$ values and fit it with Equation 3 to obtain $\kappa_{\textrm{fit}}$. In a perfect scenario, we expect $\kappa_{\textrm{fit}}$ to match the bin center value. The $\kappa_{\textrm{fit}}$ value for each bin versus the bin center value is plotted in Figure 5(b). The diagonal solid line indicates perfect agreement and the dashed lines around it show the width of a bin in $\kappa$. Looking at both Figures 5(a) and 5(b), we see good agreement in the predicted and fit $\kappa$ values in the regions where the majority of our statistics lie.
@@
Figure 6: Left: 2D histogram of predicted uncertainty ($\kappa$) and the angle from true to predicted direction for all test data. The red curve is the mean angle for each bin in $\kappa$. The black dashed curve is the expected mean angle according to Equation 3. Right: The cosine of the angles for each bin in $\kappa$ is fit to Equation 3 to obtain $\kappa_{\textrm{fit}}$, plotted with statistical error bars versus the bin center $\kappa$ value.
@@
In directional recoil detection, it is useful to break up directional performance into mean axial mismeasurement (the mean angle between the predicted axis and the true axis) and head/tail recognition efficiency (the fraction of tracks in which the head and tail ends of the recoil are predicted correctly) (Vahsen et al., 2021). This is displayed for the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing in Figure 6(a) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing in Figure 6(b). The HCN and RCN models have similar performance at 100% event efficiency, significantly outperforming the Non-ML models on both performance metrics. By omitting events with high predicted uncertainty from the HCN model, performance can be further enhanced, at the cost of event efficiency.
@@
Figure 7: Left: Head/tail recognition efficiency versus mean axial mismeasurement for all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing. The color scale indicates the event efficiency. The HCN and RCN models have similar performance at 100% event efficiency. By omitting events with high predicted uncertainty, the performance of HCN can be improved, as displayed by the curve on the top-left of the plot. Right: The same plot for the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing.

### v2

@@ 1 Introduction
In order to clarify our contribution, we summarize related work in three related areas. In Section 2.1, we review the literature on deep models for predicting direction using non-probabilistic methods or methods that do not use 3D geometry. In Section 2.2, we review the literature for the related problem of probabilistically predicting 3D orientations. In Section 2.3, we discuss the use of directional statistics for learning deep embedding spaces.
@@
Through experiments we show that our probabilistic direction predictor significantly outperforms traditional algorithms when applied to a simulated electron recoil data set. Leveraging our framework’s uncertainty predictions, performance can be enhanced by discarding high-uncertainty samples. By discarding only the top 1% of events with highest directional uncertainty, our model achieves the best expected directional performance (as defined in A). This development is far-reaching as it has the potential to enhance the sensitivity of all directional recoil detection experiments. Exciting possibilities include aiding in the discovery of dark matter (Billard et al., 2014; O’Hare, 2021) or resolving the solar abundance problem (Villante and Serenelli, 2019; O’Hare et al., 2022). The framework can also be applied to general problems where direction needs to be predicted. We demonstrate this by applying it to the task of detecting the directions 3D arrows in B.
Our contributions include: (1) The first deep learning framework that predicts 3D direction probabilistically, (2) a derivation of the negative log likelihood (NLL) loss function and approximations that stabilize training, (3) two experiments that empirically demonstrate learning and calibration, something which is lacking in related work, and (4) a deep model for predicting the direction of electron recoils from 3D point cloud data using sparse 3D convolutional networks.
@@ 2.3 Deep embedding models
Hyperspheres have long been popular manifolds for embedding data. Embedding data on the curved hypersphere provides a natural way to balance the competing attraction of similar data and repellence of dissimilar data in clustering (Banerjee et al., 2005) or multidimensional scaling (Elad et al., 2005). For the same reason, many deep learning models constrain latent space representations to a hypersphere, for example in variational autoencoders (Davidson et al., 2018) or in contrastive learning (He et al., 2020; Chen et al., 2020). Some of this work makes use of the vMF distribution to quantify the variance in heteroscedastic predictions on the hypersphere latent space (Davidson et al., 2018). However, these embedding methods are unsupervised in that the embeddings on the hypersphere are learned to optimize some objective. This work concerns the supervised learning scenario in which fixed targets live on the sphere.
@@ 3.2 Practical considerations
Some considerations are required to make the likelihood function and gradients of the probabilistic model tractable. First, the $g(\kappa_{i})\equiv\ln{\frac{\kappa_{i}}{4\pi\sinh{(\kappa_{i})}}}$ term on the right-hand side of Equation 4 is numerically unstable. When $\kappa_{i}=0$, the vMF distribution in Equation 3 is uniform over the sphere, but during learning $\kappa_{i}$ generally increases as the distribution becomes more localized (uncertainty decreases). If $\kappa_{i}\geq 87$, the denominator of $g(\kappa_{i})$ evaluates to zero (assuming a 32-bit floating point data type) resulting in an infinite loss. To avoid this, we use the limit
$\lim_{\kappa_{i}\to\infty}g(\kappa_{i})=\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}.$
In fact, $g(\kappa_{i})\approx\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}$ is a well-known approximation even for small values of $\kappa_{i}$ (Dhillon and Sra, 2003). For example, if $\kappa=9$, the values of $\ln(\frac{\kappa_{i}}{4\pi\sinh{(\kappa_{i})}})$ and $\ln(\frac{\kappa_{i}}{2\pi})-\kappa_{i}$ are indistinguishable when using the 32-bit floating point data type. Therefore, we replace the loss definition from Equation 4 with
$\textrm{NLL}=\frac{-1}{N}\sum_{i=1}^{N}\mathbbm{1}_{\kappa_{i}<9}\left(g(\kappa_{i})\right)+\mathbbm{1}_{\kappa_{\geq}9}\left(\ln\left(\frac{\kappa_{i}}{2\pi}\right)-\kappa_{i}\right)+\kappa_{i}(\mathbf{y}_{i}\cdot\hat{\mathbf{y}}_{i}).$ (5)
This function can be implemented in PyTorch (Paszke et al., 2019) using the torch.where function, but unfortunately at the time of this writing, PyTorch produces a nan gradient if either of the terms evaluates to infinity (Paszke et al., 2019). Thus we solve this problem by removing the term $g(\kappa_{i})$ from Equation 5 completely, approximating it with its 15th-order Taylor Series about $\kappa_{i}=0$
$T_{15}(\kappa_{i})\triangleq\sum^{15}_{0}\frac{g^{(n)}(0)}{n!}\kappa_{i}^{n}=\ln(4\pi)+\frac{\kappa^{2}}{6}-\frac{\kappa^{4}}{180}+\frac{\kappa^{6}}{2835}-\frac{\kappa^{8}}{37800}+\dots.$
@@
The condition $\kappa_{i}<2.65$ is chosen so that Equation 6 is continuous and is in agreement with Equation 4 within a margin of $0.14\%$.
@@ 4 Application to Electron Recoils
For a demonstration of our method on a simple toy problem see B. Here, we apply it to directional dark matter detectors. The most mature directional recoil detection technology is the gas time projection chamber (TPC). Gas TPCs are capable of determining 3D direction of recoiling atomic nuclei or electrons by reconstructing their ionization tracks. The detectors produce a 3D voxel grid, where each voxel provides a measurement of ionization at that position, as illustrated in Figure 1. We focus on determining the initial direction of the electron recoil signal because it represents the most complex scenario. There is also abundant literature on a non-machine learning algorithm for determining the direction of electrons in gas detectors (Weisskopf et al., 2016; Marco et al., 2022) which serves as a well-studied non-machine learning baseline in our experiments.
@@ 4.1 Simulation, Preprocessing, and Data splits
The final dataset contains 2,766,798 simulations. A random data split is used to create a training set (80%) and a validation set (20%). To create the test data sets we simulate another $2\times 10^{4}$ electron recoils at $40$ and 505050\,50keV. We follow the same steps outlined above except for a slight modification. The applied Gaussian smearing is no longer drawn from a uniform distribution. Instead, we probe two specific cases: a high diffusion case where $443\,{\upmu}{\rm m}$ of smearing is applied and a low diffusion case where $200\,{\upmu}{\rm m}$ is applied. The two energies and two diffusion cases give us four test data sets. Training our models across a range of energy and diffusion makes them more robust to variations in simulation parameters. Testing in specific scenarios, like high diffusion/low energy, evaluates performance in challenging conditions, while low diffusion/high energy evaluates performance in easier conditions. The specific test cases also simplify future comparisons.
@@ 4.2 Architecture
The output of the 3D convolution layers is flattened and then passed to a sequence of densely-connected layers that produce the parameters of the vMF distribution (Figure 4b). One arm ends in three linear neurons with an L2-normalization activation function, interpreted as the predicted direction on $\mathbb{S}^{2}$ ($\hat{\mathbf{y}}_{i}$), while a second arm ends with a single output neuron with a Softplus activation function, interpreted as the predicted uncertainty ($\kappa_{i}$). To evaluate the advantage of the probabilistic approach, we also train a non-probabilistic version in which the uncertainty arm is removed (Figure 4a). We refer to the probabilistic, heteroscedastic convolutional network as vMF-NN and this non-probabilistic, “deterministic” convolutional network as Det-NN. The Det-NN baseline can be viewed as the 3D generalization of the biternion neural networks introduced by Beyer et al. (2015). This represents the current state-of-the-art method for predicting direction in 3D.
Figure 4: The dense portion of our models. The feature extraction described in Figure 3 is followed by either the left or the right neural network to make the Det-NN or vMF-NN model, respectively. The illustrations do not display the actual number of neurons per layer. The number of neurons and the activation functions are specified in the legend. A learnable bias is added to every layer.
@@ 4.3 Training
The Det-NN and vMF-NN models are trained by minimising the loss functions outlined in Equations 1 and 6, respectively, using the Adam optimizer (Kingma and Ba, 2014). The learning rate is set to ${\textrm{lr}=}0.0001$, the beta values are $(0.94,0.999)$, and $\textrm{eps}=10^{-7}$, while all other parameters are kept at the PyTorch default value. The deep learning models are trained with mini-batches of size $N=256$. Hyperparameters were optimized by exploring values in the range $\textrm{eps}\in[10^{-8},10^{-6}]$, $\textrm{lr}\in[10^{-6},10^{-4}]$, and $N\in[64,256]$.
Both the Det-NN and vMF-NN models are trained on the training set while their loss on the validation set is used for early stopping and hyperparameter selection. If the validation loss has not decreased in the last five epochs, training is stopped. The weights corresponding to the lowest validation loss are saved and used for the final model.
@@ 4.4 Performance
We evaluate the trained deep learning models on the independent test sets discussed in Section 4.1 by comparing them to a commonly used non-ML algorithm and to our estimation of the best possible performance under certain idealized conditions. The non-machine learning algorithm is adapted from Marco et al. (2022) and generalized to 3D (Non-ML standard). This algorithm has three hyperparameters, so we also explore a variation (Non-ML tuned) in which the hyperparameters are fit to the test sets. The best expected directional performance is estimated by utilizing information that is only available in simulation, including the true starting point and the correct head-tail information (Best-Expected). The purpose of the Best-Expected metric is not to be a viable option for analysis but to provide a benchmark of exceptional performance. The Non-ML and Best-Expected algorithms are detailed in A.
@@
The cosine distance loss (Equation 1) is used to compare performance. Hence, when comparing the models we plot the cosine distance loss versus the percentage of omitted tracks, referred to as the efficiency cut. In Figure 5, we display the performance of all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing (left) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing (right). In both cases, we find that the deep learning models significantly outperform Non-ML, even under the most optimistic assumptions. On the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing and at a $0\%$ efficiency cut, the Det-NN, vMF-NN, and Best-Expected models achieve a cosine distance loss of $0.104$, $0.104$, and $0.098$, respectively. Converting these values to degrees by taking the arccos of $1-L_{cos}$ gives $26.4^{\circ}$, $26.4^{\circ}$, and $25.6^{\circ}$. On the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing and at a $0\%$ efficiency cut, the models achieve a cosine distance loss of $0.0624$, $0.0632$, and $0.0569$, in the same order. Performing the same conversion to degrees gives $20.3^{\circ}$, $20.5^{\circ}$, and $19.4^{\circ}$. The vMF-NN and Det-NN models have similar performance. However, the accuracy of the vMF-NN model can be improved by discarding events with low $\kappa_{i}$ (high predicted uncertainty). The vMF-NN model outperforms Best-Expected with just a $1\%$ efficiency cut, in both cases.
@@
Figure 5: Cosine distance loss versus efficiency cuts for all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing (left) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing (right). The Det-NN (red diamond) and Best-Expected (green square) models have no efficiency cuts. For the vMF-NN model, we improve accuracy by excluding examples with higher predicted uncertainty, results are presented with a blue curve. The orange point and black star indicate the performance of the Non-ML method with standard parameters and with parameters tuned on the test set, respectively. The downward arrows indicate how much improvement is possible if we help the Non-ML method by providing it with the true head-tail.
@@
To assess the calibration of the vMF-NN model we plot a 2D histogram of the predicted uncertainty ($\kappa_{i}$) and the angle from the true direction to the predicted direction ($\theta_{i}$) for all test data, displayed in Figure 6(a). The spread in $\theta_{i}$ gets smaller for higher values of $\kappa$, indicating that the vMF-NN model is predicting $\kappa_{i}$ appropriately. The solid red curve in Figure 6(a) shows the mean angle, calculated for each bin in $\kappa$. Using Equation 3, we can also calculate the expected mean angle as a function of $\kappa$
@@
where $I_{o}$ is a modified Bessel function of the first kind. The expected mean angle as a function of $\kappa$ is plotted as the dashed black line in Figure 6(a). The agreement of the expected mean angle with the mean angle for each bin in $\kappa$ suggests that the vMF-NN model is well-calibrated.
@@
Building onto the 2D histogram in Figure 6(a), we check whether the predicted angular mismeasurements in a $\kappa$ bin behave as if they are drawn from a von Mises-Fisher distribution with that value of $\kappa$. For each bin in $\kappa$, we compute a distribution of $\cos{\theta_{i}}$ and fit it with Equation 3 to obtain the maximum likelihood estimation (MLE), $\kappa_{\rm{MLE}}$. In a perfect scenario, we expect $\kappa_{\rm{MLE}}$ to match the bin center value. The $\kappa_{\rm{MLE}}$ value for each bin versus the bin center value is plotted in Figure 6(b). The diagonal solid line indicates perfect agreement and the dashed lines around it show the width of a bin in $\kappa$. Looking at both Figures 6(a) and 6(b), we see good agreement where the majority of our statistics lie. The discrepancy at higher values of predicted $\kappa$ indicates that the model is overconfident in its predictions. This is not surprising, given the known tendency for neural networks to overfit and be overconfident. However, this represents only a small fraction of the dataset.
@@
Figure 6: Left: 2D histogram of predicted uncertainty ($\kappa$) and the angle from true to predicted direction for all test data. The red curve is the mean angle for each bin in $\kappa$. The black dashed curve is the expected mean angle according to Equation 3. Right: For each bin in $\kappa$ on the left hand side, the angular mismeasurements are used to find the MLE of $\kappa$ in Equation 3. The MLE values are plotted versus the bin center $\kappa$ value.
@@
In directional recoil detection, it is useful to break up directional performance into mean axial mismeasurement (the mean angle between the predicted axis and the true axis) and head/tail recognition efficiency (the fraction of tracks in which the head and tail ends of the recoil are predicted correctly) (Vahsen et al., 2021). This is displayed for the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing in Figure 7(a) and the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing in Figure 7(b). The vMF-NN and Det-NN models have similar performance at 100% event efficiency, significantly outperforming the Non-ML models on both performance metrics. By omitting events with high predicted uncertainty from the vMF-NN model, performance can be further enhanced, at the cost of event efficiency.
@@
Figure 7: Left: Head/tail recognition efficiency versus mean axial mismeasurement for all models on the 404040\,40keV test data set with $443\,{\upmu}{\rm m}$ Gaussian smearing. The color scale indicates the event efficiency. The vMF-NN and Det-NN models have similar performance at 100% event efficiency. By omitting events with high predicted uncertainty, the performance of vMF-NN can be improved, as displayed by the curve on the top-left of the plot. Right: The same plot for the 505050\,50keV test data set with $200\,{\upmu}{\rm m}$ Gaussian smearing.